Operations Report – November 29th

MDRS Operations Report 27-29-NOV-2021

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: One of the charge controllers appears to have frozen on ABSORB setting. I noticed it yesterday and confirmed it today.

Action taken for non-nominal systems: I was not able to reset it so I will equalize it tomorrow. I didn’t want to do anything to change the system until I tried to autostart one more time and equalizing it requires that charge controller to shut down and restart.

Generator: The power system is giving me no love. For three nights I’ve set the generator to autostart and it has not. I’m giving it one more chance before I give up and run it manually. The good news is that now that the propane installation had been corrected, I can do that if I need to. I reprogrammed everything back to use SOC as the parameter for the autostart and let’s see what happens tonight. If this doesn’t work, I’m going with Occam’s Razer and assuming the problem is in the power system’s wiring, as this is the second generator that won’t autostart. But we would need to wait until spring to resolve that as our installer is out of the country. I’ve been fighting this how many years now?

ScienceDome Dual Split: Off

Solar—Nominal, supplying all power

Solar— VDC Last 24 hours:

Average 51.37

Minimum 46.50

Maximum 58.80

Since the online system is not showing SOC, I am assuming my changes to the power system did not work. Should I just bang my head against a wall now?

Notes on solar: Nothing to report

Propane Reading, station tank – 46 %

Propane Reading, director tank— 28 %

Propane Reading, intern tank— 72+ %

Propane Reading, generator— 80 %

Ethanol Free Gasoline – 1 gallons

Water (loft tank): 40 gallons

Water Meter: not checked units

Water (static tank) – 300

Static to Loft Pump used – yes

Water in GreenHab – 20 gallons (not accessible)

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 60 gallons

Hab toilet tank emptied: no

Perseverance rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Sojourner rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Curiosity rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 350.1, 350.2, 300): Nothing to report

Reason for use: n/a

Oil Added? No

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town and to quarry with media team

CrewCar used and why, where? No

Luna used and why, where? No

General notes and comments: Propane company should be here tomorrow to switch out the three Amerigas tanks.

Summary of internet: All accounts are nominal

Summary of suits and radios: Nothing to report

Campus wide inspection, if action taken, what and why? Nothing to report

Summary of general operations: Nothing to report

Summary of Hab operations: Turned on the main heater all day yesterday. No evidence of a propane leak.

Summary of Outpost operations: Nothing to report

Summary of GreenHab operations: I’ve been working in the GreenHab getting it ready for the next crew. Everything is growing well. I’ve turned on the old lights for the past three nights. They have not run as programmed. I changed the timer today to an older one and we will see what happens. We may need to order a new one.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: It’s good to see the dome open when I go out right before bed and first thing in the morning.

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Atila and Sebastian will be arriving tomorrow so this will be the last night I spend alone!

Operations Report Nov 25th – 26th

MDRS Operations Report 25-26-NOV-2021

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: n/a

Generator: I worked on the power system yesterday and today. I was getting a little desperate! Late this afternoon, it appears we are currently operational with the generator in autostart using VDC the parameter. Keep your fingers crossed! I did not change any of the power usage to make sure this actually works before heating the Hab and turning on the supplemental lighting.

ScienceDome Dual Split: Off

Solar—Nominal, supplying all power

Solar— VDC Last 24 hours (Friday):

Average 51.68

Minimum 47.18

Maximum 58.58

SOC Last 24 hours:

100 Max

68 Min

86.5 Average

Notes on solar: Nothing to report

Propane Reading, station tank – 50 %

Propane Reading, director tank— 36%

Propane Reading, intern tank— 72+ %

Propane Reading, generator— 80 %

Ethanol Free Gasoline – 1 gallons

Water (loft tank): 40 gallons

Water Meter: not checked units

Water (static tank) – 340

Static to Loft Pump used – yes

Water in GreenHab – 20 gallons (not accessible)

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 80 gallons

Hab toilet tank emptied: no

Perseverance rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Sojourner rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Curiosity rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 350.1, 350.2, 300): Nothing to report

Reason for use: n/a

Oil Added? No

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town

CrewCar used and why, where? No

Luna used and why, where? Yes, to town

General notes and comments: Nothing to report

Summary of internet: All accounts are nominal

Summary of suits and radios: Nothing to report

Campus wide inspection, if action taken, what and why? Nothing to report

Summary of general operations: Nothing to report

Summary of Hab operations: Nothing to report. Still not heating the building and the water is freezing lightly at night, then thawing during the day.

Summary of Outpost operations: Nothing to report

Summary of GreenHab operations: Nothing to report except there are too many greens ready for harvest for just one person.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Nothing at this time.

Operations Report Nov 21st – 24th

MDRS Operations Report 21-24-NOV-2021

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: n/a

Generator: Propane line has been changed to ¾ inch per manufacturer specs. In addition, since we calculated a possible use of 20 gallons a night in the middle of winter, we exchanged the 250-gallon tank for a 500-gallon tank.

ScienceDome Dual Split: Off

Solar—Nominal, supplying all power

Solar— VDC Last 24 hours (Wednesday):

Average 51.20

Minimum 46.40

Maximum 58.40

Notes on solar: Tomorrow I will reprogram the generator autostart back into the power system.

Propane Reading, station tank – 50 %

Propane Reading, director tank— 40%

Propane Reading, intern tank— 72+ %

Propane Reading, generator— 80 %

Ethanol Free Gasoline – 1 gallons

Water (loft tank): 40 gallons

Water Meter: not checked units

Water (static tank) – 370 gallons (I plugged in the heat tape as the Hab’s water lines have begun to freeze at night)

Static to Loft Pump used – yes

Water in GreenHab – 20 gallons (not accessible)

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 80 gallons

Hab toilet tank emptied: no

Perseverance rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Sojourner rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Curiosity rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 350.1, 350.2, 300): Nothing to report

Reason for use: n/a

Oil Added? No

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town

CrewCar used and why, where? Yes, to start. Started but was sluggish. Will replace battery before crew use.

Luna used and why, where? Yes, to town

General notes and comments: Nothing to report

Summary of internet: All accounts are nominal

Summary of suits and radios: Nothing to report

Campus wide inspection, if action taken, what and why? Nothing to report

Summary of general operations: Nothing to report

Summary of Hab operations: A survey of water line issues show that all three are simple fixes and will be completed soon. The toilet water line is closed and the new part arrived today. The water lines are freezing during the night and thawing by afternoon.

Summary of Outpost operations: In my trailer, a new kitchen faucet and a new bathroom were installed. The shower was fixed to operate minimally. Two 250-gallon propane tanks have been dropped off to be switched over for both trailers after Thanksgiving.

Summary of GreenHab operations: Using about 10 gallons of water a day for the plants currently growing. The LED lights arrived today and need to be installed. I’m keeping the temperature at night around 77 degrees F.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Nothing at this time.